|
ID |
Date |
Author |
Category |
Type |
Module |
Target/Number |
Subject |
Text |
 |
|
2805
|
Tuesday, May 13, 2025, 13:13 |
Frank Song | MAA | Repair | | | Swap out the target hall alpha CAM. | AlphaCAM has been replaced and tested by
RPG in MAA target hall. (per Document-154779)
WorkPermit # I2025-05-13-3 |
|
|
2804
|
Tuesday, May 13, 2025, 13:12 |
Frank Song | MAA | Repair | | | Swap out the target hall alpha CAM. |
|
|
|
2803
|
Monday, May 12, 2025, 09:17 |
David Wang | Facilities | Standard Operation | | | BL2A MAA ASU Test | MCR Travis Curucz , Frank, and I finished
BL2A MAA ASU test per document 154779. everything
works fine. Travis kept the test document. |
|
|
2802
|
Wednesday, May 07, 2025, 13:41 |
Aaron Tam | North Hot-Cell | Standard Operation | TM2 | UC#49 | UC#49 SIS w/ Gasline Install | May 7th, 2025 - Aaron Schmidt, Aaron Tam
UC#49 Handover - good |
   |
|
2801
|
Tuesday, May 06, 2025, 14:40 |
Frank Song | North Hot-Cell | Standard Operation | TM2 | UC#49 | electrical check/leak check | TM2 with new target UC#49 installed
is E/L checked on the flange of NHC in TH.
see attachment for detail. |
|
|
2800
|
Tuesday, May 06, 2025, 09:18 |
Frank Song | ITE | Standard Operation | TM4 | UCx#48 | Module Connection | TM4 with first target UC#48 of this
year has been connected in ITE. see attachment
please. |
|
|
2799
|
Monday, May 05, 2025, 09:18 |
chad fisher | North Hot-Cell | Maintenance | | | Manipulator Remote End Replacement 9352 with 9764 | Remote end 9352 replaced with end 9764
due to signs of an issue with the tong cable
(tong opening excessively large and slack |
|
|
2798
|
Friday, May 02, 2025, 10:16 |
Aaron Tam | North Hot-Cell | Maintenance | | | NHC Window Oil Fill | Findings
During April 2025 safety
walk ID: 2796 , it was noted that the |
12x |
|
2797
|
Thursday, May 01, 2025, 09:33 |
Travis Cave | ITE | Standard Operation | TM4 | UC#48 | Module move | TM#4 with UC#48 went from the conditioning
station to the east target station. |
|
|
2796
|
Friday, April 25, 2025, 14:30 |
Adam Newsome | Facilities | Standard Operation | | | Safety Walkaround complete - ISAC Hot Cells, Target Hall | A safety walkaround was completed for the
ISAC Hot Cells and Target Hall Areas.
The resulting spreadsheet can be |
|
|
2795
|
Thursday, April 24, 2025, 13:42 |
David Wang | Conditioning Station | Standard Operation | | | Refilled TCS:TK2 from 11" to 28.8 " | Refilled TCS:TK2 from 11" to 28.8
". |
|
|
2794
|
Thursday, April 24, 2025, 12:53 |
Aaron Tam | North Hot-Cell | Standard Operation | TM4 | UC#48 | UC#48 IGLIS Install onto TM4 | C. Fisher, A. Schmidt, A. Tam
UC#48 Handover satisfactory
Target taken out of glovebox |
 |
|
2793
|
Thursday, April 24, 2025, 11:33 |
Travis Cave | Conditioning Station | Standard Operation | | UC#48 | Module move | TM4 with UC48 went on a long trip
from the north hot cell to the conditioning
station. |
|
|
2792
|
Thursday, April 24, 2025, 11:13 |
Frank Song | North Hot-Cell | Standard Operation | TM4 | UCx#48 | electrical check/leak check | TM4 with new target UCx#48 installed
was E/L checked on the top of NHC, see attachment
uploaded with pls. |
|
|
2791
|
Wednesday, April 23, 2025, 14:08 |
Adam Newsome | North Hot-Cell | Standard Operation | TM4 | | TM4 Module Move | TM4 was moved from Safe Module Parking
to the North Hot Cell today. |
|
|
2790
|
Friday, April 11, 2025, 13:46 |
chad fisher | North Hot-Cell | Repair | | | Remote end 9**4 repair | Tong cable on remote arm 9**4 has be relocated
onto its pulleys and tong opening adjusted
at the tong cable drum by changing the position |
|
|
2789
|
Friday, April 11, 2025, 13:34 |
chad fisher | North Hot-Cell | Repair | | | remote end 9342 installed | On March 27 remote end 9342 was installed
on the west position (right hand as viewed
by operator) and arm 9764 removed and moved |
|
|
2788
|
Friday, April 04, 2025, 09:35 |
David Wang | Crane | Development | | | spent target pail and its lifting jigs load test | Adam and I did spent target pail
and its lifting jigs load test this morning.
The load for full pail test is 263 lbs and |
|
|
2787
|
Tuesday, April 01, 2025, 09:47 |
Frank Song | North Hot-Cell | Repair | | | Two NHC light bulbs replacement | 2 broken lights of NHC were replaced in
Target Hall. see pics attached. |
    |
|
2786
|
Thursday, March 20, 2025, 13:10 |
David Wang | Conditioning Station | Standard Operation | | | TCS water system filters annual replacement has been done. | TCS water system filters annual replacement
has been done. Plant group Dan Tower finished
the job and updated the inspection tags . |
|